### v1

# VersaPRM: Multi-Domain Process Reward Model via Synthetic Reasoning Data

Abstract
Process Reward Models (PRMs) have proven effective at enhancing mathematical reasoning for Large Language Models (LLMs) by leveraging increased inference-time computation. However, they are predominantly trained on mathematical data and their generalizability to non-mathematical domains has not been rigorously studied. In response, this work first shows that current PRMs have poor performance in other domains. To address this limitation, we introduce VersaPRM, a multi-domain PRM trained on synthetic reasoning data generated using our novel data generation and annotation method. VersaPRM achieves consistent performance gains across diverse domains. For instance, in the MMLU-Pro category of Law, VersaPRM via weighted majority voting, achieves a 7.9% performance gain over the majority voting baseline—surpassing Qwen2.5-Math-PRM’s gain of 1.3%. We further contribute to the community by open-sourcing all data, code and models for VersaPRM.
Process Reward Models, Multi-Domain Process Reward Models, Synthetic Reasoning Data
1 Introduction
Large Language Models (LLMs) have demonstrated significant potential in tackling complex reasoning tasks. Specifically, they can employ a step-by-step Chain-of-Thought (CoT) approach to generate more accurate and reliable solutions (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023; Madaan et al., 2023). Moreover, by using additional test-time computation, the reasoning performance of LLMs can be further enhanced (Snell et al., 2024; Yao et al., 2024).
An important and widely-adopted test-time computation method is using external verifiers, such as reward models to rank multiple generated solutions and select the best answer (Lightman et al., 2024). Reward models evaluate the quality of solutions, helping guide LLMs toward better outputs. In particular, Outcome Reward Models (ORMs) are used to provide supervision based solely on the correctness of the final outcome. However, ORMs fail to address errors in intermediate steps, limiting their effectiveness for complex, multi-step reasoning tasks (Luo et al., 2024; Lightman et al., 2024; Sun et al., 2024). Because ORMs suffer from this limitation, Process Reward Models (PRMs) have been proposed to offer fine-grained, step-by-step feedback on the correctness of each reasoning step (Lightman et al., 2024; Uesato et al., 2022). PRMs have proven highly effective during inference, improving the reranking of generated solutions and guiding LLMs through search-based algorithms (Wan et al., 2024; Wang et al., 2024a).
Figure 1: Existing open-source PRMs trained on math datasets achieve strong math performance and can outperform a majority voting baseline when used via weighted majority voting. However, these PRMs fail to generalize to other domains (e.g., Law, Philosophy, and Biology), performing no better than the baseline. We propose a multi-domain PRM, VersaPRM, obtained by further fine-tuning a math PRM on a synthetically generated multi-domain dataset. The resulting PRM effectively generalizes beyond math, improving test-time reasoning across multiple domains.
Several studies have shown that PRMs trained on extensive process supervision significantly outperform ORMs in mathematical reasoning tasks, with notable improvements reported on datasets such as MATH500 and GSM800K (Luo et al., 2024; Lightman et al., 2024; Uesato et al., 2022). While substantial investigation has been made in creating training data (Lightman et al., 2024; Wang et al., 2024b), training PRMs (Xiong et al., 2024), and evaluation (Zheng et al., 2024) with respect to mathematical reasoning, the application of PRMs to non-mathematical domains—such as Biology, Chemistry, and Law—remains underexplored. To investigate the capability of math PRMs in non-mathematical domains, we test open-source math PRMs such as Math-Shepherd (Wang et al., 2024b) and Qwen-2.5-Math-PRM (Zheng et al., 2024). Not surprisingly, these PRMs demonstrate poor performance, indicating their limited domain generalizability. They exhibit only marginal improvements over the baseline in Law, Philosophy, and Biology as illustrated in Figure 1.
To address this limitation, we propose fine-tuning PRMs on a synthetically generated multi-domain CoT dataset, to significantly enhance reasoning capabilities beyond mathematics. We call this resulting multi-domain PRM VersaPRM, short for versatile PRM. Notably, by sampling questions from the MMLU-Pro dataset (Wang et al., 2024c), we generate CoTs to produce step-by-step reasoning using an LLM-based generator, i.e., Llama-3.1-8B-Instruct (Dubey et al., 2024), and then auto-label them using an LLM-based labeler, i.e., Llama-3.1-70B-Instruct (Dubey et al., 2024). VersaPRM, which is trained on the resulting synthetic multi-domain reasoning data, shows strong performance across diverse domains. We validate its superior performance using various empirical analyses on VersaPRM against existing open-source PRMs.
Our contributions are summarized as follows:
  1. 1.
    We identify the limited domain generalizability of open-source math PRMs in Section 4.
  2. 2.
    We propose a novel data generation and annotation pipeline across multiple domains in Section 5.
  3. 3.
    We introduce a large-scale, high-quality multi-domain process supervision dataset, dubbed MMLU-Pro-CoT-Train (Labeled).
  4. 4.
    We train a well-generalized PRM that outperforms existing baselines, demonstrating strong generalization across diverse domains in Section 6.
  5. 5.
    We open-source the implementation of VersaPRM with training details, our multi-domain reasoning data, and its model checkpoint; available at https://github.com/UW-Madison-Lee-Lab/VersaPRM.
2 Related Work
Outcome Reward and Process Reward Models.
PRMs have proven more effective than ORMs in enhancing LLM reasoning, particularly for mathematical tasks (Luo et al., 2024; Lightman et al., 2024; Sun et al., 2024). Unlike ORMs, which focus on final outcomes, PRMs provide step-by-step feedback, improving error detection in intermediate steps and multi-step tasks (Luo et al., 2024; Lightman et al., 2024; Uesato et al., 2022; Wang et al., 2024b). Techniques like OmegaPRM (Luo et al., 2024) and Math-Shepherd (Wang et al., 2024b) reduce reliance on costly human annotations, while RLHflow (Xiong et al., 2024), OpenR (Wang et al., 2024a) and ProcessBench (Zheng et al., 2024) advance PRM evaluation and training. However, the expertise of existing PRMs is mainly limited to mathematical reasoning. Our work extends the capability of PRM to multi-domain reasoning, using synthetic reasoning data for broader applicability.
Test-Time Inference Algorithms.
Test-time inference algorithms enhance LLM reasoning by adding computation during inference. AlphaCode (Li et al., 2022) shows how test-time computing boosts competitive programming performance, while Snell et al. (2024) argue that scaling inference-time compute, rather than model parameters, yields better results by adapting compute allocation to prompt difficulty. Test-time inference includes self-improvement methods like Tree of Thoughts (ToT) (Yao et al., 2024), self-verification (Weng et al., 2023), and stepwise self-evaluation (Xie et al., 2024), and external-verifier methods such as verifier reranking (Cobbe et al., 2021; Wan et al., 2024), tool feedback (Gou et al., 2024), reward-guided reasoning (Yang et al., 2024b), and multi-agent debate (Du et al., 2024). Our work uses PRMs for reranking solutions and guiding reasoning within the external verifier paradigm.
Synthetic Data Generation.
Obtaining fine-grained step-wise labeling of CoTs via expert annotation is costly and time-consuming. Automated annotation methods, such as rollout-based approaches, reduce human effort but require numerous model inference, which is computationally expensive. Recent effort to mitigate these limitations include multi-rollout generation per reasoning step (Wang et al., 2024b, d) and efficiency improvements via binary search (Luo et al., 2024). Despite these advancements, the reliance on multiple model calls remains a bottleneck. Recent studies by Gilardi et al. (2023) and Fonseca & Cohen (2024) demonstrate the potential of LLMs as data labelers. In line with the work of Zheng et al. (2024), our work uses LLMs as labelers to automatically generate process reward annotations, enabling the production of synthetic data across multiple domains without high computational costs.
3 Process Reward Models
Similar to the work of Lightman et al. (2024), we define process rewards to represent the correctness of each step, and whether it is logical and follows from previous steps.
To formally define a PRM, we begin by specifying a CoT $S=(s_{1},s_{2},\dots,s_{k})$ as a sequence of $k$ reasoning steps, where $s_{i}$ is the $i$-th step in the CoT for each $i\in[k]$. A PRM can then be formally characterized as a function that maps each CoT $S$ to an associated $k$-dimensional vector of rewards: $\text{PRM}(S)\in[0,1]^{k}$. The $i$-th coordinate of the output score vector, denoted as $\text{PRM}(S)_{i}$, represents the PRM score for the correctness of the reasoning step $s_{i}$.
3.1 Score Aggregation Methods
Using a PRM, we can obtain scores for each reasoning step. To then scalarize the reward score vector of the whole CoT, we consider the following three aggregation methods.
Min-Aggregation. We use the minimum PRM step score in a CoT as the aggregated score:
$\text{Aggr}_{\text{min}}(S)=\min_{i\in[k]}\text{PRM}(S)_{i}.$
Last-Aggregation. We utilize the PRM score of the last step in a CoT as the aggregated score:
$\text{Aggr}_{\text{last}}(S)=\text{PRM}(S)_{k}.$
Average-Aggregation. We employ the average PRM step score of the CoT as the aggregated score:
$\text{Aggr}_{\text{avg}}(S)=\frac{1}{k}\sum_{i\in[k]}\text{PRM}(S)_{i}.$
These aggregated scores are particularly useful for solution reranking and are employed in the test-time inference algorithms described below (Wang et al., 2024b; Sun et al., 2024; Lightman et al., 2024).
3.2 Inference-Time Methods
In this section, we introduce three reranking-based methods—Majority Voting, Weighted Majority Voting, and Best-of-$N$—along with two search-based methods—Beam Search and Monte Carlo Tree Search.
Let $a_{S}$ denote the final answer in a CoT $S$, which in practice can be extracted using a suitable parser. Further let $\mathcal{S}_{N}=\{S_{1},S_{2},\dots,S_{N}\}$ denote a set of $N$ CoTs sampled i.i.d. from a generator over a particular question.
Majority Voting (MV). MV (Wang et al., 2023) is a robust baseline inference-time method that does not require a PRM. Specifically, we first sample $N$ candidate solutions to a problem from a generator. The final answer is then determined by selecting the solution that appears most frequently among these $N$ candidates:
$\text{MV}(\mathcal{S}_{N})=\operatorname*{arg\,max}_{a_{S}:S\in\mathcal{S}_{N}}\sum_{i\in[N]}\mathds{1}_{a_{S}}(a_{S_{i}}).$
Weighted Majority Voting (WMV). This method, as used by Uesato et al. (2022), is similar to MV. We still first sample $N$ candidate solutions. However, we weight the frequencies of CoTs with identical answers by the aggregation scores. The final answer is the one with highest sum of weights:
$\text{WMV}(\mathcal{S}_{N})=\operatorname*{arg\,max}_{a_{S}:S\in\mathcal{S}_{N}}\sum_{i\in[N]}\mathds{1}_{a_{S}}(a_{S_{i}})\cdot\text{Aggr}(S_{i}).$
Best-of-$N$ (BoN). This method also samples $N$ candidate solutions. It then reranks them using the aggregation score from a PRM. The answer of the solution with highest score is chosen as final answer:
$\text{BoN}(\mathcal{S}_{N})=\operatorname*{arg\,max}_{a_{S}:S\in\mathcal{S}_{N}}\text{Aggr}(S).$
Beam Search. This method (Snell et al., 2024) is initialized with a fixed number of beams $N$ and width $M$. The process starts by sampling $N$ initial predictions for the first reasoning step. These are ranked via the PRM’s step score, retaining the top $\frac{N}{M}$ candidates. For each retained candidate, $M$ proposals for the next step are sampled, yielding $N$ new candidates. This iterates until all beams reach solutions or a maximum iteration limit. The final prediction is selected based on the highest aggregated PRM score across steps. See Algorithm 1 for details.
Monte Carlo Tree Search (MCTS). MCTS is a search algorithm used during test-time inference (Hao et al., 2023; Wan et al., 2024) that iteratively builds a search tree to find the CoT with the highest aggregated PRM score. A detailed description is presented in Appendix B and the pseudo-code is provided in Algorithm 2.
4 Limitations of Process Reward Models Trained on Math Domain Data
We introduce various math PRMs used for comparison in Section 4.1, present our multi-domain evaluation dataset in Section 4.2, and provide a detailed analysis of the evaluation results in Section 4.3.
4.1 Open-Source Math PRMs
For evaluation, we conduct experiments on a diverse set of models. Our analysis includes four open-source math PRMs: Math-PSA (Wang et al., 2024a), Math-Shepherd (Wang et al., 2024b), RLHFLow-Deepseek (Xiong et al., 2024), and Qwen-2.5-Math-PRM (Zheng et al., 2024).
In addition to the open-source models, two math PRMs based on open-source models are specifically trained in this work. They are denoted as LlamaPRM800K and QwenPRM800K. More details are given in Section C.1
4.2 Multi-Domain Evaluation Dataset
For our multi-domain evaluation dataset, we curate questions sampled from the MMLU-Pro dataset (Wang et al., 2024c). MMLU-Pro is designed to benchmark the reasoning abilities of LLMs and consists of college-level multiple choice questions in the following 14 domains: Math, Physics, Chemistry, Law, Engineering, Other, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, and History.
To craft our evaluation dataset, we randomly sample 150 questions from each domain. Due to duplicate questions, we discard 41 questions—23 from Biology, 10 from Health, 5 from Law, and 1 each from Business, Economics, and Philosophy. For each remaining question, we generate 128 candidate solutions using Llama-3.1-8B-Instruct (Dubey et al., 2024) for MV, WMV, and BoN test-time inference algorithms. Prompt details and generation parameters are provided in Section A.2. We refer to this multi-domain evaluation dataset as MMLU-Pro-CoT-Eval (Unlabeled).
Table 1: Results of two open-source math PRMs on different domains in MMLU-Pro-CoT-Eval (Unlabeled) when using WMV with min-aggregation on 16 CoTs generated per question using Llama-3.1-8B-Instruct. In parenthesis we report absolute difference between WMV and MV (WMV--MV). While WMV using math PRMs exhibits greater improvement in Math and Math-adjacent domains, there is no significant improvement on MV in other domains.
Category MV Math-Shepherd Qwen-2.5-Math-PRM
All 57.15 57.66 (+0.51) 58.17 (+1.02)
All except math 56.61 57.01 (+0.40) 57.32 (+0.71)
Math 62.40 64.13 (+1.73) 67.20 (+4.80)
Chemistry 58.67 60.13 (+1.46) 60.67 (+2.00)
Physics 58.53 61.87 (+3.34) 61.47 (+2.94)
Biology 75.38 75.38 (+0.00) 75.69 (+0.31)
Psychology 61.60 61.47 (-0.13) 62.27 (+0.67)
Law 35.93 37.24 (+1.31) 36.28 (+0.35)
History 49.20 49.87 (+0.67) 49.40 (+0.20)
Philosophy 44.83 44.70 (-0.13) 45.17 (+0.34)
4.3 Multi-Domain Performance of Math PRMs
We conduct comprehensive analyses on a diverse set of models. For clarity, we report results for two representative models here, with additional evaluations available in Section D.1. The first model, Math-Shepherd (Wang et al., 2024b), is trained on synthetically generated math data labeled via a rollout-based method. The second model, Qwen-2.5-Math-PRM (Zheng et al., 2024), is a best-performing open-source PRM, trained on the high-quality expert labeled PRM800K math dataset (Lightman et al., 2024).
The PRMs are applied using WMV with min-aggregation. While math PRMs show significant improvements in mathematical reasoning domains, their effectiveness in broader, non-mathematical areas remains limited. Notably, in the Math category, Qwen-2.5-Math-PRM and Math-Shepherd achieve relative gains of $+4.80$ and $+1.73$, respectively, outperforming the MV baseline. Similar improvements are observed in Math-adjacent disciplines: Chemistry ($+2.00$ for Qwen-2.5-Math-PRM) and Physics ($+3.34$ for Math-Shepherd), underscoring their utility in tasks requiring mathematical reasoning.
Finding 1:
Math PRMs struggle to generalize to broader domains.
However, the benefits diminish sharply in non-mathematical areas. For example, in Philosophy and History, we see gains of only $+0.34$ and +0.20% respectively for the most performant PRM Qwen-2.5-Math-PRM.
The “All except math” aggregate further underscores this disparity, with PRMs achieving a maximum gain of $+0.71$ (Qwen-2.5-Math-PRM) compared with the majority voting baseline.
These results highlight a critical limitation: math PRMs trained exclusively on mathematical data lack the versatility to generalize beyond mathematical reasoning tasks. While they excel in contexts aligned with their training—quantitative reasoning—their capacity to evaluate reasoning quality in broader domains remains insufficient.
5 Automatic Generation of Multi-Domain Reasoning Data with Labels
Figure 2: A diagram of the synthetic data generation pipeline. In the CoT Generation Stage, each question is used to generate 16 CoT solutions. Then, in the Auto-Labeling Stage, each CoT is evaluated to create step-wise labels. If a CoT step is labeled as BAD, all subsequent steps will be discarded.
In order to obtain step-wise reasoning data for non-Math domains, we devise a pipeline, as outlined in Figure 2, to generate synthetic reasoning CoTs from existing question-answering data. These CoTs are then given step-wise labels based on reasoning correctness. We detail the synthetic data generation process in Sections 5.1 and 5.2, including methods to create and annotate reasoning steps. We also provide additional analysis on the quality of the generation pipeline in Section 5.3.
5.1 Chain-of-Thought Generation
For the generation of CoTs, we prompt Llama-3.1-8B-Instruct to produce step-by-step reasoning for each input question. For training, we source questions from the MMLU-Pro dataset (Wang et al., 2024c), selected for its high-quality, challenging problems spanning diverse topics. From this dataset, we randomly sample up to 500 questions per domain, ensuring that it is disjoint to the subset used for evaluation. We then generate 16 CoTs for each sampled question. Post-generation, we filter out CoTs exceeding the 2,048-token limit or containing unparsable answers.
5.2 Auto-Labeling
To annotate our synthetic CoT data, we adopt an approach inspired by the critic models in the work of Zheng et al. (2024). Specifically, we utilize Llama-3.1-70B-Instruct as a strong LLM to evaluate each CoT using step-by-step reasoning, locating the earliest erroneous step, if any. To enhance accuracy and consistency, we identified two key additional components.
First, we incorporate explicit step evaluation definitions, inspired by Lightman et al. (2024), into the system prompt. Steps are categorized as GOOD, OK, or BAD: BAD for incorrect, unverifiable, or irrelevant steps; GOOD for correct, verifiable, and well-aligned steps; OK for intermediate cases. Second, we also provide the reference ground-truth answer in the prompt. The full prompt is detailed in Section A.2.
To convert the auto-labeling output to stepwise labels, we apply the following rule: if no steps are detected as incorrect, all steps in the CoT are labeled as $1$. If a step is detected as incorrect, all preceding steps are labeled as $1$, the incorrect step is labeled as $-1$, and all subsequent steps are discarded.
In total, we sample 5,750 questions from MMLU-Pro. Among the 84,098 generated CoTs that passed filtering, 36,935 were labeled as having no incorrect steps and 47,163 were labeled as having at least one (see Table 2). This dataset, denoted as MMLU-Pro-CoT-Train (Labeled), is the first open-source multi-domain reasoning dataset with step-wise labels.
To assess the quality of our auto-labeled data, we conduct a manual evaluation on a random sample of 30 questions from the dataset. For each question, we randomly select one CoT classified as entirely correct and two CoTs flagged as containing an incorrect step. We then manually validate whether the auto-labeled judgments align with our own assessments.
For the CoTs labeled as correct by the auto-labeler, we observed an agreement rate of 83% with our manual evaluations. For CoTs labeled as incorrect, the agreement rate was 70%.
Based on these results, we estimate that approximately 75% of the CoTs in the entire dataset are correctly labeled. This level of accuracy is comparable to that of manually-labeled CoT datasets, such as PRM800K (Lightman et al., 2024), which is estimated to achieve around 80% accuracy. [^1] Refer to this GitHub issue for a discussion on PRM800K’s accuracy.
5.3 Auto-Labeling Prompt Analysis
To further understand the factors influencing auto-labeling performance, we conduct an evaluation of the auto-labeling using a simplified prompt. Specifically, we remove the system prompt defining the types of reasoning steps and exclude the reference ground-truth answer from the prompt. When re-evaluating the auto-labeling quality, we observed a drastic drop in performance, with the agreement rate for CoTs labeled as correct by the original auto-labeler decreasing by over 70%, from 83% to 7%, while the agreement rate for CoTs labeled as incorrect decreased from 70% to 62%.
These results highlight the importance of providing both a well-defined prompt with step label definitions and access to the ground-truth answer in achieving high auto-labeling accuracy. The ground-truth answer provides essential context on CoT final correctness and enhances the model’s ability to evaluate reasoning steps effectively.
5.4 Counterfactual Augmentation
To generate additional examples of incorrect reasoning, we explore methods for instructing an LLM to modify steps in our correct CoTs, introducing specific types of errors. We refer to this process broadly as counterfactual augmentation. However, incorporating counterfactual error steps during PRM training was not observed to significantly improve performance. Therefore, we defer specific details and experiments using counterfactual augmentation to Section A.3.
6 Multi-Domain Process Reward Model
We present the implementation and evaluation of VersaPRM, structured as follows. First, Section 6.1 covers the various training configurations used. We then evaluate VersaPRMvia BoN and WMV in Section 6.2, showing improved domain generalization compared to math PRMs. In Section 6.3, we additionally discuss results using Beam Search and MCTS. Lastly, we examine VersaPRM’s ability to scale test-time compute for larger models such as Deepseek-R1 (Guo et al., 2025) in Section 6.4.
6.1 Training of Our Multi-Domain PRM
To train VersaPRM, we employ a classification head atop an LLM, optimizing with a cross-entropy loss applied to a special classification token appended at the end of each CoT step in MMLU-Pro-CoT-Train (Labeled). Detailed specifics and hyperparameters are provided in Section C.2.
We explore several training configurations, including: 1) LoRA (Hu et al., 2022) vs. full fine-tuning for efficient training, 2) a base LLM vs. a math PRM for initializing the PRM, and 3) a Qwen-based PRM vs. a Llama-based PRM for training. Comprehensive experimental results for these studies are presented in the next section. Based on those findings, our final, our final multi-domain PRM, named VersaPRM, is initialized from our LlamaPRM800K—see Appendix C for its details—fine-tuned using LoRA on our multi-domain training dataset.
6.2 Math PRM vs. VersaPRM on Reranking Based Inference-Time Methods
We first report results of the reranking methods WMV and BoN on MMLU-Pro-CoT-Eval (Unlabeled). For both methods, we adopt Min-aggregation, as it outperforms Average and Last in aggregating PRM step scores; see Section D.4 for comparison. We also include MV as a baseline.
Comparison with Math Open-Source PRMs. We evaluate our multi-domain PRM, VersaPRM, against open-source math PRMs by partitioning MMLU-Pro-CoT-Eval (Unlabeled) into three groups: 1) Math, 2) Math-adjacent, i.e., Chemistry, Computer Science, Engineering, Physics, and 3) non-Math-adjacent domains. As shown in Figure 3, our model consistently outperforms baselines in both WMV and BoN across all domain groups.
Finding 2:
Fine-tuning with synthetic multi-domain data enhances the generalizability of PRM.
For WMV, we can see the relative performance difference increase with domain distance from core mathematics. While performance of open-source math PRMs converges to the majority voting baseline in non-mathematical domains, our multi-domain PRM maintains robust generalization.
In BoN the superiority of our multi-domain PRM is even more pronounced. Unlike open-source math PRMs, which fail to surpass the baseline of MV in Math-adjacent and non-Math-adjacent domains, our model consistently surpasses it across all domain groups.
See Section D.6 for more fine-grained details where we plot WMV and BoN for every domain of MMLU-Pro-CoT-Eval (Unlabeled). The results are consistent with Figure 3, and VersaPRM outperforms math PRMs in all domains.
Figure 3: Comparison of WMV (top) and BoN (bottom) using VersaPRM against open-source math PRMs on MMLU-Pro-CoT-Eval (Unlabeled). We use min-aggregation and the CoTs are generated using Llama-3.1-8B-Instruct. VersaPRM has consistently better performance than math PRMs, and the differences become larger in domains not adjacent to Math.
Ablation Experiments Using Multi-Domain PRM Trained on Math Only Subset vs. Random Subset.
We further conduct an ablation study to evaluate the impact of training data diversity on the performance of our LlamaPRM800K Math PRM. Specifically, we train one PRM using only the math subset of our multi-domain training data and another using a random subset of the same size. We refer to these two models as VersaPRM (Math subset) and VersaPRM (random subset), respectively. This experiment tests that the improved performance of our multi-domain PRM is due to the domain-diversity of the CoT data and not merely from learning the in-distribution question and CoT formats of MMLU-Pro questions. If the latter is the case, both PRMs should perform similarly, given that they are exposed to the same amount of questions and CoT examples with the in-distribution format.
Finding 3:
Domain diversity of CoTs in a training dataset plays an integral role in generalization of PRMs to multiple domains.
As shown in Figure 4, VersaPRM (random subset) achieves superior performance in WMV compared to VersaPRM (Math subset). This trend holds across both Math and non-Math domains. These findings suggest two key insights. First, our PRM is not simply learning the question format but is acquiring knowledge on how to label reasoning across diverse domains. This is why training on diverse data enables better overall performance than training on same sized data in only one domain. Second, VersaPRM (random subset) also demonstrates slightly better performance in the math domain, indicating that training on a diverse dataset may facilitate positive transfer, where insights from other domains enhance reasoning in the Math domain.
Figure 4: Comparison of WMV using LlamaPRM800K, VersaPRM (Math subset) and VersaPRM (random subset). VersaPRM (random subset) achieves better performance than VersaPRM (Math subset) in Math and non-Math.
Experiments Using Other Training Configurations. While our final version of VersaPRM is trained from LlamaPRM800K on our synthetic data using LoRA, we also test the following training configurations on our multi-domain dataset:
  • VersaPRM (Llama Base): We initialize training from Llama-3.1-8B-Instruct, and use LoRA fine-tuning with our multi-domain dataset.
  • VersaPRM (Qwen): We initialize training from QwenPRM800K PRM, and utilize LoRA fine-tuning with our multi-domain dataset.
  • VersaPRM (full-tuned): We initialize training from LlamaPRM800K PRM, and do full fine-tuning with our multi-domain dataset.
The results are presented in Figure 5. Comparing VersaPRM (Qwen) and VersaPRM (Llama), we observe that the QwenPRM800K VersaPRM performs worse. This highlights the importance of base model choices. Although Qwen-2.5-Math-7B, the base model for QwenPRM800K, is specialized in mathematical reasoning, its limitations in general-domain knowledge hinder its ability to fully leverage multi-domain training data.
Finding 4:
Exposure to mathematical data beforehand can enhance a PRMs’ ability to effectively leverage multi-domain CoT fine-tuning.
Next, comparing VersaPRM (Llama Base) with VersaPRM, we find that the latter achieves superior performance in Math while maintaining comparable performance in non-Math domains. This suggests that prior exposure to mathematical data enhances the model’s ability to benefit from further domain-specific training.
We note that VersaPRM (full-tuned) has worse performance than VersaPRM. This may be due to suboptimal hyperparameters leading to overfitting during full fine-tuning.
Figure 5: Comparison of MVW using VersaPRM against other multi-domain PRMs trained using different configurations. VersaPRM has better WMV performance than all other models in both Math and non-Math domains.
6.3 Math PRM vs. Multi-Domain PRM on Search Based Inference-Time Methods
We evaluate the performance of math PRMs (using LlamaPRM800K) and VersaPRM with beam search and MCTS on MMLU-Pro-CoT-Eval (Unlabeled). The results over questions in all domains, presented in Figure 6, show that MCTS outperforms beam search and that they both do better than the MV baseline. Regardless of the search algorithm used, consistent with our WMN and BoN results, VersaPRM gives boosted performance over the math PRM. Details by category results are presented in Section D.7.
Figure 6: Comparison of VersaPRM and LlamaPRM800K with beam search and MCTS. Overall in the diverse domains from MMLU-Pro-CoT-Eval (Unlabeled), VersaPRM achieves better performance.
6.4 Does PRM with Test-Time Compute help Reasoning Models?
Figure 7: Comparison of WMV using VersaPRM against Qwen-2.5-Math-PRM on DeepSeek-R1 generated CoTs for the Law subset. VersaPRM has better performance than all other math PRMs.
We have shown that VersaPRM can effectively leverage inference-time compute to increase LLM performance, a natural question is whether this effectiveness extends to renowned strong reasoning models, e.g., DeepSeek-R1 (Guo et al., 2025). Given that a well-trained reasoning model may already generate coherent and correct reasoning steps due to being trained for reasoning, one might hypothesize that reranking methods like WMV and BoN brings marginal improvement over WM.
To test this, we evaluate the performance of VersaPRM via WMV on DeepSeek-R1. Due to budget constraints, we focus on the Law subset and sample 16 CoT responses per question. As shown in Figure 7, VersaPRM provides a slight but noticeable performance boost to DeepSeek-R1 during test-time inference despite the limited CoT samples. Significantly it outperforms both the math PRM and the MV baseline. This finding, though preliminary, nullifies the aforementioned hypothesis and suggests that—in fact—large reasoning models can still benefit from PRMs during inference to further boost their performance beyond MV.
7 Discussion and Future Directions
We proposed VersaPRM trained using synthetic reasoning data to address the limitations of existing math PRMs. By leveraging a cost-efficient synthetic data generation pipeline, we enabled production of high-quality step-wise reasoning data and demonstrate that PRMs can effectively scale reasoning ability at inference time in diverse domains.
Future Work. Several directions remain for advancing multi-domain PRMs. Can VersaPRM be effectively used as a reward model for RL training? Can it improves RL training beyond math PRMs? Could more sophisticated counterfactual augmentation enhance PRM effectiveness? Also, evaluating PRMs on harder, open-ended problems would better assess their generalization. Lastly, more thorough exploration of PRMs with large models (e.g., GPT-4 (Achiam et al., 2023) and DeepSeek-R1) could clarify their scalability and role in state-of-the-art reasoning systems.
Acknowledgements
Kangwook Lee is supported by NSF CAREER Award CCF-2339978, an Amazon Research Award, and a grant from FuriosaAI. In addition, Thomas Zeng acknowledges support from NSF under NSF Award DMS-202323 and Daewon Chae was supported by the Hyundai Motor Chung Mong-Koo Foundation.
Impact Statement
Given the potential for LLMs to be used in unethical ways, such as spreading misinformation or manipulating public opinion, our Multi-Domain PRM could inadvertently contribute to such misuse. To mitigate these risks, it is essential to implement robust safeguards in training and inference.
Appendix A More Details on Synthetic Data Generation Pipeline
A.1 Dataset Composition
The total composition of MMLU-Pro-CoT-Train (Labeled) is as follows.
Table 2: Composition of MMLU-Pro-CoT-Train (Labeled)
Total Fully Correct Incorrect
Number of CoTs 84098 36935 47163
Number of Steps 487380 440217 47163
A.2 Data Generation Pipeline Prompts
To generate chain-of-thought (CoT) reasoning for MMLU-Pro questions, we utilize the prompt shown in Figure 11. To ensure the generated CoT adhere to the proper format—where steps are separated by two newline characters and the final step follows the structure “the answer is (X)”—we include five few-shot examples. These examples are derived from the CoTs provided in the validation split of MMLU-Pro, with additional processing to ensure each step is delimited. The code for generating the complete prompt will be open-sourced alongside the rest of our code and data.
During generation, we use a temperature of $0.8$ and set the maximum generation length to 2,048 tokens. During auto-labeling, we use a temperature of 0, and the maximum generation length remains at 2,048 tokens.
The following is a multiple choice question and its ground truth answer. You are also given a students solution (split into step, enclosed with tags and indexed from 0): [Multiple Choice Question] {question} [Ground Truth Answer] {answer} [Student Solution] {<<<step_0>>> Student solution step 0 <<</step_0>>> <<<step_1>>> Student solution step 0 <<</step_1>>> …}
Figure 8: User prompt template for auto-labeling.
[…]
Figure 9: System prompt for auto-labeling.
A.3 Counterfactual Augmentation
Figure 10: Diagram of the counterfactual augmentation pipeline
After generating and labeling our synthetic reasoning CoTs (as described in Section 5), we attempted to create additional incorrect steps by augmenting the correct reasoning steps. Our pipeline is depicted in Figure 10. We provide the full CoT to Llama-3.1-70B-Instruct, instructing it to select and rewrite a step where it would be appropriate to introduce an error. Additionally, we define a list of possible fine-grained error types. To encourage the generation of a variety of different error types, we only include a random selection of two of these error types in each system prompt, forcing the LLM to choose one. The error types are:
  • Conflicting Steps: The reasoning step includes statements that contradict previous steps.
  • Non-sequitur: The reasoning step introduces information that is irrelevant to the question.
  • Factual: The reasoning step contains incorrect statements of factual information.
  • False Assumption: The reasoning step makes an incorrect assumption about the question.
  • Contextual: The reasoning step misinterprets information given within the question/context.
For the prompt format used in counterfactual augmentation, see Figure 12. In total, we generated 73,829 augmented incorrect steps.
The following are multiple choice questions (with answers). Think step by step and then finish your answer with ”the answer is (X)” where X is the correct letter choice.
Figure 11: Prompt to generate CoTs for MMLU Pro.
You are a highly knowledgeable philosopher with expertise across many domains, tasked with analyzing reasoning processes. Your goal is to identify how a reasoning process could naturally deviate toward an incorrect conclusion through the introduction of subtle errors. Here are a list of potential error types, all of which are equally valid:
[ERROR TYPE 1]: [ERROR TYPE 1 DEFINITION]
[ERROR TYPE 2]: [ERROR TYPE 2 DEFINITION]
Instructions:
You will be provided with:
1. A question.
2. A complete chain of reasoning steps, where each step is numbered (e.g., Step X).
Your task is to: 1. Identify the major factual information, reasoning, and conclusions within the reasoning steps.
3. Explain how to generate an incorrect step to replace one of the existing steps. This should include:
- Identifying a step where the reasoning could naturally deviate.
- Speculating what type of error would be most appropriate to introduce at the chosen step.
4. Introduce an incorrect next step that aligns stylistically with the previous steps. This incorrect step should:
- Reflect a deviation in reasoning that significantly harms the correctness.
- Appear natural and believable in the context of the reasoning process.
5. Clearly explain how the incorrect step is an error, highlighting the specific logical or conceptual flaw.
Output Format: STEP_SUMMARY:
[Summarize the reasoning within the steps in 1-2 sentences, identifying major information, logical steps, and conclusions.]
INCORRECT_STEP_GEN:
[Explain how the reasoning at a specific step could deviate naturally into being incorrect. Clearly describe the type of error that could be introduced at this step.]
ERROR_TYPE:
[The name of the type of error chosen to be introduced.]
STEP_NUM:
[The number of the step that was identified as a place where the reasoning could naturally deviate. Only include the number here.]
INCORRECT_STEP:
[Write the incorrect step in the same tone and style as the other steps. Wrap the incorrect step inside curly braces (e.g. {incorrect step}).]
ERROR_EXPLANATION:
[Explain how the incorrect step fits the definition of the selected error type, identifying the specific flaw.]
Figure 12: System prompt for counterfactual augmentation.
Appendix B Additional Search Algorithm Details
Algorithm 1 Beam Search with Process Reward Model
0:  Large Language Model $\text{LLM}(\cdot)$, Process Reward Model $\text{PRM}(\cdot)$, Prompt $s_{0}$, Number of Beams $N$, Beam width $M$, Maximum step length $L$
1:  $\mathcal{B}\leftarrow[s_{0}]$
2:  $\mathcal{Q}\leftarrow[0]$
3:  for $i=1$ to $L$ do
4:     $\mathcal{B}\leftarrow\text{Expand}(\mathcal{B},\frac{N}{\operatorname{len}(\mathcal{B})})$
5:     $\mathcal{B}\leftarrow\text{LLM.step}(\mathcal{B})$
6:     $\mathcal{Q}\leftarrow\text{Aggr}(\mathcal{B})$
[…]
10:     if All sequences in $\mathcal{B}$ contain a terminal leaf node then
11:        break
12:     end if
13:  end for
14:  Return the sequence with the highest score from $\mathcal{B}$
Algorithm 1 is a greedy search algorithm that uses a PRM select the best CoT during search. More details are given in Section 3.2.
Algorithm 2 Monte Carlo Tree Search with Process Reward Model
0:  Large Language Model $\text{LLM}(\cdot)$, Process Reward Model $\text{PRM}(\cdot)$, Prompt $s_{0}$, Maximum step length $L$, Number of roll-outs $K$, Number of generated child nodes $d$, Exploration weight $w$
1:  Initialize the value function $Q:\mathcal{S}\mapsto\mathbb{R}$ and visit counter $N:\mathcal{S}\mapsto\mathbb{N}$
2:  for $n\leftarrow 0,\dots,K-1$ do
3:     // Selection
4:     $t\leftarrow 0$
5:     while $s_{t}$ is not a leaf node do
6:        $N(s_{t})\leftarrow N(s_{t})+1$
7:        $s_{t+1}\leftarrow\arg\max_{\text{children}(s_{t})}\left[Q(\text{child}(s_{t}))+w\sqrt{\frac{\ln N(s_{t})}{N(\text{child}(s_{t}))}}\right]$
8:        $t\leftarrow t+1$
9:     end while
10:     // Expansion & Simulation (equivalent to the beam search with $N=M=d$)
11:     $\mathcal{B}\leftarrow[s_{t}]$
12:     while $s_{t}$ is not a terminal leaf node $\wedge$ $t\leq L$ do
13:        $N(s_{t})\leftarrow N(s_{t})+1$
14:        $\mathcal{B}\leftarrow\text{Expand}(\mathcal{B},d)$
15:        $\mathcal{B}\leftarrow\text{LLM.step}(\mathcal{B})$
16:        for $s\in\mathcal{B}$ do
17:           $Q(s)\leftarrow\text{Aggr}(s)$
18:        end for
19:        $s_{t+1}\leftarrow\arg\max_{s\in\mathcal{B}}Q(s)$
20:        $t\leftarrow t+1$
21:        $\mathcal{B}\leftarrow[s_{t}]$
22:     end while
23:     // Back Propagation
24:     for $t^{\prime}\leftarrow t,\dots,0$ do
25:        $Q(s_{t^{\prime}})\leftarrow\max(Q(s_{t^{\prime}}),Q(s_{t}))$
26:     end for
27:  end for
28:  Return the sequence with the highest score among the terminal nodes
Algorithm 2 is a tree-based search algorithm that iteratively expands a search tree to find the CoT with the highest PRM score. MCTS iteratively builds a search tree through the following steps:
  1. 1.
    Selection: Starting from the root node, the algorithm traverses the tree by selecting child nodes according to a selection policy.
  2. 2.
    Expansion and Simulation: Upon reaching a non-terminal leaf node, the tree is expanded iteratively by generating a fixed number of child nodes and then greedily selecting the child node with the highest value (which for us is determined by the PRM). This process continues until a terminal node is reached.
  3. 3.
    Backpropagation: The results from the simulation are propagated back through the tree, updating value estimates and visit counts for each node along the path.
These steps are repeated for a fixed number of iterations or until a computational or time limit is reached. To determine the final prediction, we choose the terminal node with the highest value.
Appendix C Additional PRM Training Details
C.1 Open-Source Math PRM Training Details
The open-source PRMs evaluated in this work utilize CoT training data derived from two mathematical datasets: MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021). The Math-Shepherd and RLHFlow/Deepseek-PRM-Data datasets are synthetically generated following the rollout method proposed by Wang et al. (2024b). Similarly, the MATH-APS dataset is produced using the synthetic generation technique introduced by Luo et al. (2024). PRM800K, in contrast, consists of manually annotated labels and was specifically curated for the study by Lightman et al. (2024).
All PRMs are trained using the base LLMs of comparable model size and class, including Mistral-7B (Jiang et al., 2023), Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen-2.5-Math 8B (Yang et al., 2024a).
Table 3: Training details of various Math PRMs
PRM Base Model Training Data Training Method
Math-PSA Qwen-2.5-Math-7B-Instruct PRM800K, Math-Shepherd and MATH-APS LoRA
Math-Shepherd Mistral-7B Math-Shepherd Full fine-tuning
Qwen-2.5-Math-PRM Qwen-2.5-Math-7B-Instruct PRM800K Full fine-tuning
RLHFLow-Deepseek Llama3.1-8B-Instruct RLHFlow/Deepseek-PRM-Data Full fine-tuning
LlamaPRM800K Llama3.1-8B-Instruct PRM800K Full fine-tuning
QwenPRM800K Qwen-2.5-Math-7B-Instruct PRM800K Full fine-tuning
C.2 Details of PRM Training
[…]
For training our math PRMs on the PRM800K dataset (QwenPRM800K and LlamaPRM800K), we employ a batch size of 128 and perform full fine-tuning. For experiments on mixed-domain datasets, we reduce the batch size to 32 due to smaller dataset size.
All training is conducted over a single epoch. For full fine-tuning, we use a learning rate of $1.25\times 10^{-6}$, while for LoRA-based fine-tuning, we use a learning rate of $1.0\times 10^{-4}$.
Appendix D Additional PRM Training and Evaluation Experiments
D.1 Evaluation Results for Math PRMs and VersaPRM Across all Categories
Table 4: Comparison among various math PRMs and VersaPRM on different domains in MMLU-Pro-CoT-Eval (Unlabeled) when using WMV with min-aggregation on $N=16$ CoTs generated per question using Llama3.1-8B-Instruct. In parenthesis we report the relative difference between WMV and the MV baseline (WMV--MV). While WMV using math PRMs exhibit greater improvement in math and math-adjacent domains, there is no significant improvement on MV in other domains.
Category MV (Baseline) Math-PSA Math-Shepherd Qwen-2.5-Math-PRM RLHFLow-Deepseek LlamaPRM800K VersaPRM
All 57.15 57.87(+0.72) 57.66(+0.51) 58.17(+1.02) 57.59(+0.44) 58.16(+1.01) 61.22(+4.07)
All except math 56.61 56.82(+0.21) 57.01(+0.40) 57.32(+0.71) 56.96(+0.35) 57.71(+1.10) 60.29(+3.68)
Math 62.40 64.20(+1.80) 64.13(+1.73) 67.20(+4.80) 64.07(+1.67) 65.40(+3.00) 68.87(+6.47)
Math-Adjacent 56.75 57.98(+1.23) 57.48(+0.73) 58.30(+1.55) 57.33(+0.58) 58.27(+1.52) 61.22(+4.47)
Non-Math-Adjacent 56.69 56.79(+0.10) 57.14(+0.45) 57.09(+0.40) 57.02(+0.33) 57.55(+0.86) 60.00(+3.31)
Chemistry 58.67 60.47(+1.80) 60.13(+1.46) 60.67(+2.00) 59.13(+0.46) 60.47(+1.80) 66.13(+7.46)
Computer Science 55.80 56.93(+1.13) 56.07(+0.27) 56.13(+0.33) 56.07(+0.27) 56.40(+0.60) 58.60(+2.80)
Engineering 51.67 50.67(-1.00) 51.07(-0.60) 53.13(+1.46) 51.87(+0.20) 52.27(+0.60) 55.27(+3.60)
Physics 58.53 61.87(+3.34) 61.87(+3.34) 61.47(+2.94) 60.80(+2.27) 61.47(+2.94) 64.87(+6.34)
Biology 75.38 75.23(-0.15) 75.38(+0.00) 75.69(+0.31) 75.77(+0.39) 76.38(+1.00) 80.00(+4.62)
Health 63.36 63.00(-0.36) 63.93(+0.57) 63.50(+0.14) 63.57(+0.21) 64.50(+1.14) 65.50(+2.14)
Psychology 61.60 61.47(-0.13) 61.47(-0.13) 62.27(+0.67) 61.47(-0.13) 61.87(+0.27) 64.53(+2.93)
Business 61.34 61.95(+0.61) 62.21(+0.87) 63.02(+1.68) 62.21(+0.87) 62.62(+1.28) 64.50(+3.16)
Economics 62.00 62.67(+0.67) 62.33(+0.33) 62.53(+0.53) 62.67(+0.67) 62.40(+0.40) 64.27(+2.27)
Law 35.93 35.72(-0.21) 37.24(+1.31) 36.28(+0.35) 36.07(+0.14) 36.90(+0.97) 43.86(+7.93)
History 49.20 49.00(-0.20) 49.87(+0.67) 49.40(+0.20) 49.40(+0.20) 49.87(+0.67) 50.67(+1.47)
Philosophy 44.83 44.90(+0.07) 44.70(-0.13) 45.17(+0.34) 44.56(-0.27) 45.30(+0.47) 49.13(+4.30)
Other 55.53 55.80(+0.27) 55.47(-0.06) 56.07(+0.54) 55.87(+0.34) 57.07(+1.54) 59.00(+3.47)
D.2 PRM Training with Counterfactual Augmented Data
Figure 13: Comparison of WMV (top) and BoN (bottom) using our two multi-domain PRMs (w/ and w/o counterfactually augmented training data) on the categories of MMLU-Pro-CoT-Eval (Unlabeled). We use min-aggregation and the CoTs are generated using Llama-3.1-8B-Instruct. When using WMV, counterfactual augmented data can further improve the performance of PRM on non-math-adjacent domains.
D.3 WMV and BoN using different aggregation methods
Figure 14: Comparison of WMV (left) and BoN (right) using VersaPRM with different reward aggregations on MMLU-Pro-CoT-Eval (Unlabeled). The CoTs are generated using Llama 3.1 8B Instruct. Overall, min-aggregation brings the largest inference performance boost.
D.4 Larger Generator Inference with PRM Rewarding
Figure 15: Comparison of WMV (left) and BoN (right) using VersaPRM against math PRMs on MMLU-Pro-CoT-Eval (Unlabeled). We use min-aggregation and the CoTs are generated using Llama-3.1-70B-Instruct. Similar trends to using 8B model as the generator are observed, indicating that our Multi-Domain PRM can generalize across generators with different capacities.
D.5 Inference with Compact PRM
Figure 16: Comparison of WMV (top) and BoN (bottom) using VersaPRM (Llama3B Base), a compact PRM based on Llama-3.2-3B-Instruct and trained on our multi-domain dataset. We use min-aggregation and the CoTs are generated using Llama-3.1-8B-Instruct. Compared with using VersaPRM (Llama Base), which applies the same training data and configurations but is based on Llama-3.1-8B-Instruct, VersaPRM (Llama3B Base) brings a less significant performance boost. However, the overall trends are similar, indicating the efficacy of the inference pipeline using PRM.
D.6 Comparison of VersaPRM against Other Open-Source Math PRMs on WMV and BoN by Category
Figure 17: Comparison of WMV using VersaPRM against open-source PRMs on more other categories of MMLU-Pro-CoT-Eval (Unlabeled). We use min-aggregation and the CoTs are generated using Llama-3.1-8B-Instruct.
Figure 18: Comparison of BoN using VersaPRM against open-source PRMs on more other categories of MMLU-Pro-CoT-Eval (Unlabeled). We use min-aggregation and the CoTs are generated using Llama-3.1-8B-Instruct.
D.7 Comparison of VersaPRM against Other Open-Source Math PRMs on MCTS and Beam Search by Category
Figure 19: Comparison of VersaPRM and LlamaPRM800K with beam search and MCTS. In more other categories from MMLU-Pro-CoT-Eval (Unlabeled), VersaPRM achieves better performance.